OpenStudy (anonymous):

Subtract these polynominals. (6x^3-4x+5)-(3x^3-5x^2+6x-2) A.) 3x^3+5x^2+2x+7 B.) 3x^3-5x^2x+7 C.) 3x^3-5x^2-10x+7 D.) 3x^3+5x^2-10x+7

Vocaloid (vocaloid):

combine like terms

Vocaloid (vocaloid):

6x^3 - 3x^3 = ?

OpenStudy (anonymous):

What would be the answer ?

Vocaloid (vocaloid):

hint: 6 - 3 = 3 6x^3 - 3x^3 = ?

OpenStudy (anonymous):

Ok what would be the next step ?

Vocaloid (vocaloid):

I'm asking you to solve the expression

Vocaloid (vocaloid):

6x^3 - 3x^3 = __x^3 fill in the blank

OpenStudy (anonymous):

3